OpenStudy (anonymous):

AwakeReality Suppose the sum of three consecutive even integers is given. How can you find the three numbers?

OpenStudy (anonymous):

Well, call the middle integer k. Then the other integers are k-2 and k+2. Their sum is \[(k-2)+k+(k+2) = 3k \]and say you know what that is. So \(3k = n\). Then \(k=\frac{3}{n}\). In other words, to find the middle number, divide the sum (which you know) by 3. Then to get the other two integers, add 2 or take away 2.

OpenStudy (anonymous):

Correction: \(k = \frac{n}{3}\). That was an unfortunate typo...

OpenStudy (anonymous):

Thank you!